May 20-23

– Four stores provided ~47.25 hours of luminosity
– NuMI off for target repairs
– Tevatron personnel conducted hollow-beam studies
– Tevatron quenched during transfer of antiprotons
– EE support repaired Tevatron quadrupole regulator’s load cable
– Lightning strike on the evening of May 22 caused recycler kickers to prefire, the loss of the stash, and lost communication with Pelletron

*The integrated luminosity for the period from 5/16/11 to 5/23/11 was 59.22 inverse picobarns. NuMI took 4.73E16 of beam during this same time period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts